please help me if any of you can please..........
if m+p+k=70,p=2m and k=2p,then m=?

Answer by ewatrrr(24785)

 
Hi
p=2m and k=2p
m + p + k=70 Substitute for k
m + p + 2p = 70 Substitute for p
m + 2m + 2(2m) = 70
7m = 70
m = 10